# Secure Outsourced Decryption for HE-based Privacy-preserving Cloud Computing System

### v1

@@ Abstract
The demand for processing vast volumes of data has surged dramatically due to the advancement of machine learning technology. Large-scale data processing necessitates substantial computational resources, prompting individuals and enterprises to turn to cloud services. Accompanying this trend is a growing concern regarding data leakage and misuse. Homomorphic encryption (HE) is one solution for safeguarding data privacy, enabling encrypted data to be processed securely in the cloud. However, we observe that encryption and decryption routines of some HE schemes require considerable computational resources, presenting non-trivial work for clients. In this paper, we propose an outsourced decryption protocol for RLWE-based HE schemes, which splits the original decryption into two routines, with the computationally intensive part executed remotely by the cloud. Its security relies on an invariant of the NTRU-search problem with a newly designed secret distribution. Cryptographic analyses are conducted to configure protocol parameters across varying security levels. Our experiments demonstrate that the proposed protocol achieves up to a $67\%$ acceleration in the client’s local decryption, accompanied by a $50\%$ reduction in space usage.
@@
privacy-preserving computation, outsourced computing, homomorphic encryption, lattice-based cryptography
@@ I Introduction
Homomorphic encryption is a promising technique capable of addressing these concerns and safeguarding data privacy. Its development has seen rapid strides over the past 15 years, particularly since Gentry’s seminal breakthrough [1] in 2009. Presently, the most feature-rich and efficient homomorphic encryption schemes [2, 3, 4, 5] can facilitate secure computation of complex machine learning operations, such as neural network inference and training [6, 7, 8, 9]. These schemes are built based on the RLWE (Ring-Learning with Error) security assumption [10, 11], which relates to difficult lattice-based problems and may provide potential resilience against the attacks of quantum algorithms.
A classic example of secure cloud computing utilizing homomorphic encryption is illustrated in Figure 1. In this scenario, the client transmits its homomorphically encrypted data to the cloud. The cloud then processes the encrypted data according to the client’s function $f$ and returns the encrypted result to the client, which can be decrypted using the secret key $sk_{A}$. Despite the delegation of task $f$ to the cloud server, clients have to frequently perform encryption and decryption operations to upload data and recover results. For many RLWE-based HE schemes, these routines involve polynomial multiplications with large degrees and modulus, posing a nontrivial challenge for resource-constrained devices.
@@
To address this challenge, we propose a novel outsourced decryption protocol to ease clients’ computational burden. This protocol divides the original decryption algorithm into two stages, which can be performed collaboratively by both the cloud server and the client. The stage executed by the cloud is referred to as ’partial decryption’, while the stage executed by the client is called ’local decryption’. The latter is more efficient and requires less memory than the original decryption process. Our contributions are outlined as follows:
  • We devise a multiplication-communicative secret key blinding technique as the core of the outsourced decryption protocol. This allows us to delegate part of the decryption computation to the cloud using a blinded private key. The proposed blinding method utilizes the inverses of sparse polynomials as blinding factors, enabling unblinding to be accomplished more swiftly than executing dense polynomial multiplication, which constitutes the primary complexity of the decryption algorithm.
@@
  • We implement the protocol in the renowned homomorphic encryption library Lattigo and compare its performance with the original decryption algorithm, which shows the efficiency of our protocol.
@@ II-A Homomorphic Encryption
One of the most prevailing and versatile types of homomorphic encryption [2, 5, 4, 3] is built based on the RLWE security assumption [14, 10, 11]. It includes HE schemes that support homomorphic operations in various plaintext spaces, including complex vector space[2, 15], modulo integer vector space[16, 4, 5], $\{0,1\}$ binary space[17, 3], and others. From an algebraic standpoint, the performance of these RLWE-based HE schemes is determined by the efficiency of polynomial ring element operations, among which polynomial multiplication may be the most important factor. The Number Theoretic Transformation(NTT) is now widely deployed in HE schemes for polynomial multiplication acceleration[15, 18], boasting an asymptotic complexity of $O(d\log{d})$, where $d$ signifies the polynomial degree. Nonetheless, polynomial multiplication remains a non-trivial computation in HE schemes since $d$ may range from $2^{10}$ to $2^{16}$ along with the length of the coefficient modulus spanning from tens to thousands [19]. Typically, encryptions and decryptions in RLWE-based HE schemes require polynomial multiplications, so it is worth thinking about whether there is any way to reduce their computational burden for clients in some cloud computing scenarios. Our work may provide some interesting insight on this topic.
In the following, we briefly introduce the RLWE-based public key encryption (PKE) [14, 20, 5, 2], which is the foundation of many RLWE-based HE schemes, where routines supporting homomorphic computation are straightly built upon this PKE system (but we do not go this far since we only focus on decryption).
@@ III Protocol Overview & Security Model
The proposed outsourced decryption protocol (abbreviated as OtsDec) consists of a tuple of routines, they are $\Pi_{\text{Setup}}$, $\Pi_{\text{SkBdKeyGen}}$, $\Pi_{\text{SkBd}}$, $\Pi_{\text{ParDec}}$, $\Pi_{\text{LocDec}}$, $\Pi_{\text{CtBd}}$, and $\Pi_{\text{Verify}}$:
  • $\Pi_{\text{Setup}}$(Setup): Takes parameters $(d,q)$ as input to setup the protocol. $d$ and $q$ denote the polynomial ring $R_{q}$ used in OtsDec, which are typically inherited from the underlying homomorphic encryption scheme. The following routines take $d$ and $q$ as input implicitly.
  • $\Pi_{\text{SkBdKeyGen}}$(Blinding key generation for secret key): Outputs a pair of factors $t^{-1},t\in R_{q}$ as the key pair for blinding and unblinding the secret key.
  • $\Pi_{\text{SkBd}}$(Secret Key Blinding): Takes a secret key $sk$ from the underlying homomorphic encryption and the blinding key pair $(t^{-1},t)$ as input, and outputs the blinded secret key $\tilde{s}$.
  • $\Pi_{\text{ParDec}}$(Partial Decryption): Takes a ciphertext $ct=(u,v)\in R_{q}^{2}$ and the blinded secret key $\tilde{s}$ as input, and outputs the partial decryption $\tilde{m}$ of $ct$.
  • $\Pi_{\text{LocDec}}$(Local Decryption): Takes a partially decrypted ciphertext $\tilde{m}$, perform unblinding on it and outputs a plaintext $m$.
@@
Outsourced Decryption Phase Input: $\mathcal{T}$ inputs $\perp$; $\mathcal{U}_{H}$ inputs a set of ciphertexts $\{ct_{i}\}$ encrypted under $s$, which is typically the result of some homomorphic evaluation and to be sent to $\mathcal{T}$ for decryption. Output: $\mathcal{T}$ receives the decryption result of $\{ct_{i}\}$; $\mathcal{U}_{H}$ receives $\perp$. 1. [Partial Decryption]: $\mathcal{U}_{H}$ initializes an empty set $P$. For each ciphertext $ct=(u,v)\in R_{q}^{2}$ in $\{ct_{i}\}$, $\mathcal{U}_{H}$ performs $\prod_{\text{ParDec}}(\tilde{s},ct)$ and adds the output into $P$. $\mathcal{U}_{H}$ then sents $P$ to $\mathcal{T}$. 2. [Local Decryption]: On receiving partially decrypted results $P$, for each $\tilde{pt}\in P$, $\mathcal{T}$ performs $\prod_{\text{LocDec}}(t,\tilde{pt})$ and accepts the output as the decryption result.
@@
As shown in Figure 2, the client $\mathcal{T}$ blinds his or her secret key and sends it to the cloud server during the Setup phase. Subsequently, whenever $\mathcal{U}_{H}$ finishes a computation task ordered by $\mathcal{T}$, it performs partial decryption on the encrypted results and sends it back to $\mathcal{T}$ for local decryption.
Similar to many privacy-preserving computation schemes using HE, we employ the semi-honest security model for the proposed protocol, wherein the server $\mathcal{U}_{H}$ is assumed to be a passive adversary that adheres to the protocol but still maintains curiosity about $\mathcal{T}$’s private data.
Some existing frameworks account for malicious behaviors from $U_{H}$ and require verifiable outsourced computations. However, this seems to hold little significance for our scenario, as verifying the correctness of homomorphic operations is itself a non-trivial task and remains one of the major research goals today. Nevertheless, we will discuss feasible approaches to correctness verification at the end of the paper for future research reference.
We adopt the following notations from [29] to measure the proposed protocol’s security, correctness, and efficiency.
For our outsourcing decryption protocol to be applicable to the aforementioned scenario, it must adhere to the following properties:
@@
Correctness For all arithmetic functions $f:\mathcal{M}^{I}\rightarrow\mathcal{M}$ defined in the underlying HE scheme, there exists an instance of the outsourcing decryption protocol such that $\Pi_{LocDec}(\Pi_{ParDec}(\text{HE.Eval}(f,pk,ct_{1},\dots,ct_{I}),\tilde{sk}),sk_{ob})=f(m_{1},\dots,m_{I})$ holds with overwhelming probability.
@@ IV-A Secret Key Blinding from Sparse Polynomial Inversion
From the description above, one can observe that routines $\Pi_{\text{SkBd}}$ and $\Pi_{\text{ParDec}}$ are just normal polynomial multiplications (see Algorithms 2 and 3). Note that the former one is invoked only once in the protocol, whose computational complexity can be amortized across multiple subsequent ciphertext decryption operations, making it negligible in the overall computational load. The latter is typically performed by $\mathcal{U}_{H}$ for each ciphertext to be decrypted.
The remaining challenge is to generate blinding factors that ensure $\Pi_{\text{LocDec}}$ operates with maximal efficiency. Note that the sparse-to-dense polynomial multiplication exhibits lower complexity compared to NTT-based dense multiplication when the sparsity and the circuit design are appropriately handled. Therefore, we choose sparse polynomials as the unblinding factors in $\Pi_{SkBdKeyGen}$. The sampling procedure is outlined in Algorithm 1, where we use hamming weight (denoted as $h$) to control the sparsity of a polynomial, and $S_{h}$ the set of all sparse polynomials over $R_{q}$ with hamming weight $h$. The unblinding factor is sampled from $S_{h}$ uniformly and randomly, with its inverse being the blinding factor. Note that an invertibility test should be done for the candidate unblinding factor since $R_{q}$ is merely a ring rather than a field.
@@
Elements individually satisfying the invertibility condition from different $R_{q_{i}}$ can be combined into an invertible element in $R_{q}$ using the CRT map (see Theorem IV.1). It can also be proven that elements sampled this way are uniformly distributed across $S_{h}$ (see Theorem IV.2). Using CRT can significantly reduce the complexity of the invertibility check since $q$ may be very large as we mentioned previously. It is worth mentioning that in homomorphic encryption schemes, the structure $q=\prod_{i=0}^{L}q_{i}$ is typically used to construct a large $q$, where $q_{i}$ are primes with similar length and satisfy the condition $q_{i}\equiv 1\mod{2d}$ to enable NTT-based polynomial multiplication.
Algorithm 1 OtsDec.SkBdKeyGen (Blinding key generation for secret key)
0:  Ring $R_{q}=\mathbb{Z}_{q}[X]/X^{d}+1$; Sparsity $h$
0:  Deobfuscation factor $t$ with sparsity $h$; Obfuscation factor $t^{-1}\in R_{q}$
@@ IV-B Local Decryption
Algorithm 2 OtsDec.SkBd (Secret Key Blinding)
0:  Obfuscation factor $r\in R_{q}$; Secret key $s\in R_{q}$
0:  Obfuscated secret key $\tilde{s}\leftarrow rs\in R_{q}$
Algorithm 3 OtsDec.ParDec (Partial Decryption)
0:  Obfuscated secret key $\tilde{s}$; Ciphertext $c=(u,v)\in R_{q}^{2}$
0:  Obfuscated decryption result $\tilde{c}\leftarrow(u\tilde{s},v)\in R_{q}^{2}$
Algorithm 4 OtsDec.LocDec (Local Decryption)
0:  Deobfuscation factor $t\in R_{q}$ with sparsity $h$; Obfuscated decryption result $\tilde{c}=(\tilde{u},v)\in R_{q}$;
0:  Deobfuscated decryption result $m\in R_{q}$;
@@ IV-C Complexity Analysis
$\mathcal{T}$ is responsible for routine $\Pi_{SkBdKeyGen},\Pi_{SkBd}$ and $\Pi_{LocDec}$ in the protocol, where the former two are only performed once following the generation of the secret key, and the later is done for each ciphertext partially decrypted by $\Pi_{ParDec}$. Therefore, the complexity of $\mathcal{T}$ is dominated by the sparse-to-dense polynomial multiplication with roughly $hd$ (modular) scalar multiplications.
On the other hand, the original decryption computed by $\mathcal{T}$ has a typical complexity of $ad\log{d}+d$ scalar multiplications corresponding to one Inversed NTT operation and coordinate-wise multiplication, where $a$ is a constant varies from concrete implementations of NTT. Then the efficiency of our proposed protocol is computed as $\alpha=\frac{h}{a\log{d}+1}$.

### v2

# Secure Outsourced Decryption for FHE-based Privacy-preserving Cloud Computing

@@ Abstract
The demand for processing vast volumes of data has surged dramatically due to the advancement of machine learning technology. Large-scale data processing necessitates substantial computational resources, prompting individuals and enterprises to turn to cloud services. Accompanying this trend is a growing concern regarding data leakage and misuse. Homomorphic encryption (HE) is one solution for safeguarding data privacy, enabling encrypted data to be processed securely in the cloud. However, the encryption and decryption routines of some HE schemes require considerable computational resources, presenting non-trivial work for clients. In this paper, we propose an outsourced decryption protocol for the prevailing RLWE-based fully homomorphic encryption schemes. The protocol splits the original decryption into two routines, with the computationally intensive part executed remotely by the cloud. Its security relies on an invariant of the NTRU-search problem with a newly designed blinding key distribution. Cryptographic analyses are conducted to configure protocol parameters across varying security levels. Our experiments demonstrate that the proposed protocol achieves up to a $67\%$ acceleration in the client’s local decryption, accompanied by a $50\%$ reduction in space usage.
@@
Privacy-preserving computation, Outsourced computing, Homomorphic encryption
@@ I Introduction
Homomorphic encryption is a promising technique capable of addressing these concerns and safeguarding data privacy. Fully homomorphic encryption (FHE), in particular, supports arbitrary computation on ciphertexts. Its development has seen rapid strides over the past 15 years[1]. Presently, the prevailing FHE schemes [2, 3, 4, 5] can facilitate secure computation of complex machine learning operations, such as neural network inference and training [6, 7, 8, 9]. These schemes are built based on the RLWE (Ring-Learning with Error) security assumption [10, 11], which relates to difficult lattice-based problems and may provide potential resilience against the attacks of quantum algorithms.
A classic example of secure cloud computing utilizing homomorphic encryption is illustrated in Figure 1. In this scenario, the client transmits its homomorphically encrypted data to the cloud. The cloud then processes the encrypted data according to the client’s instruction $f$ and returns the encrypted result to the client, which can be decrypted using the secret key $sk_{A}$. Despite the delegation of task $f$ to the cloud server, clients have to frequently perform encryption and decryption operations to upload data and recover results. For RLWE-based HE schemes mentioned above, these routines involve polynomial multiplications with large degrees and modulus, posing a nontrivial challenge for resource-constrained devices.
@@
To ease clients’ computational burden, we propose a new protocol for outsourcing the decryption routine. It divides the original decryption algorithm into two stages that can be performed collaboratively by both the cloud server and the client. The stage executed by the cloud is referred to as ’blind decryption’, while the stage executed by the client is called ’local decryption’. The latter is more efficient and requires less memory than the original decryption process. Our contributions are outlined as follows:
  • We devise a multiplication-communicative secret key blinding technique as the core of the outsourced decryption protocol. This allows us to securely delegate part of the decryption computation to the cloud using a blinded secret key. The proposed blinding method utilizes the inverses of sparse polynomials as blinding factors, enabling unblinding to be accomplished more swiftly than executing dense polynomial multiplication, which constitutes the primary complexity of the decryption algorithm.
@@
  • We implement the protocol in the renowned homomorphic encryption library Lattigo[14] and compare its performance with the original decryption algorithm, which shows the efficiency of our protocol.
@@ II-A Homomorphic Encryption
One of the most prevailing and versatile types of (fully) homomorphic encryption [2, 5, 4, 3] is built based on the RLWE security assumption [15, 10, 11]. It includes HE schemes that support homomorphic operations in various plaintext spaces, including complex vector space[2, 16], modulo integer vector space[17, 4, 5], $\{0,1\}$ binary space[18, 3], and others. The performance of these RLWE-based HE schemes is determined by the efficiency of polynomial ring element operations, among which polynomial multiplication may be the most important factor. The Number Theoretic Transformation(NTT) is now widely deployed in HE schemes for polynomial multiplication acceleration[16, 19], boasting an asymptotic complexity of $O(d\log{d})$, where $d$ signifies the polynomial degree. Nonetheless, polynomial multiplication remains a non-trivial computation in HE schemes since $d$ may range from $2^{10}$ to $2^{16}$ along with the length of the coefficient modulus spanning from tens to thousands [20]. Typically, encryptions and decryptions in RLWE-based HE schemes require polynomial multiplications, so it is worth thinking about whether there is any way to reduce their computational burden for clients in some cloud computing scenarios. Our work may provide some interesting insight on this topic.
In the following, we briefly introduce the RLWE-based public key encryption (PKE) [15, 21, 5, 2], which is the foundation of RLWE-based HE schemes, where routines supporting homomorphic computation are straightly built upon this PKE system (but we do not go this far since we only focus on decryption).
@@ III Protocol Overview & Security Model
The proposed outsourced decryption protocol consists of a tuple of routines, they are $\Pi_{\text{Setup}}$, $\Pi_{\text{SkBdKeyGen}}$, $\Pi_{\text{SkBd}}$, $\Pi_{\text{BldDec}}$, and $\Pi_{\text{LocDec}}$:
  • $\Pi_{\text{Setup}}$(Setup): Takes parameters $(d,q,\chi_{sk},\lambda)$ from the underlying HE scheme as input and outputs a hamming weight $h$, where $d$ and $q$ denote the polynomial ring $R_{q}=\mathbb{Z}[X]_{q}/(X^{d}+1)$, $\chi_{sk}$ denotes the secret key distribution, and $\lambda$ denotes the security parameter.
  • $\Pi_{\text{SkBdKeyGen}}$(Blinding key generation for secret key): Takes a hamming weight $h$ as input and outputs a pair of factors $t^{-1},t\in R_{q}$ as the key pair for blinding and unblinding the secret key.
  • $\Pi_{\text{SkBd}}$(Secret Key Blinding): Takes a secret key $sk$ from the underlying HE scheme and the blinding key pair $(t^{-1},t)$ as input, and outputs the blinded secret key $\tilde{s}$.
  • $\Pi_{\text{BldDec}}$(Blind Decryption): Takes a ciphertext $c=(u,v)\in R_{q}^{2}$ and the blinded secret key $\tilde{s}$ as input, and outputs the blind decryption of $c$.
  • $\Pi_{\text{LocDec}}$(Local Decryption): Takes a blindly decrypted ciphertext $\tilde{c}$, perform unblinding on it and outputs a plaintext $m$.
@@
Outsourced Decryption Phase Input: $\mathcal{T}$ inputs $\perp$; $\mathcal{U}_{H}$ inputs a set of ciphertexts $\{ct_{i}\}$ encrypted under $s$, which is typically the result of some homomorphic evaluation and to be sent to $\mathcal{T}$ for decryption. Output: $\mathcal{T}$ receives the decryption result of $\{ct_{i}\}$; $\mathcal{U}_{H}$ receives $\perp$. 1. [Blind Decryption]: $\mathcal{U}_{H}$ initializes an empty set $P$. For each ciphertext $ct=(u,v)\in R_{q}^{2}$ in $\{ct_{i}\}$, $\mathcal{U}_{H}$ performs $\prod_{\text{BldDec}}(\tilde{s},ct)$ and adds the output into $P$. $\mathcal{U}_{H}$ then sents $P$ to $\mathcal{T}$. 2. [Local Decryption]: On receiving blindly decrypted results $P$, for each $\tilde{pt}\in P$, $\mathcal{T}$ performs $\prod_{\text{LocDec}}(t,\tilde{pt})$ and accepts the output as the decryption result.
@@
As shown in Figure 2, the proposed protocol has two phases, Setup and Outsourced decryption. During the Setup phase, the client $\mathcal{T}$ blinds his or her secret key and sends it to the cloud server. Subsequently, whenever $\mathcal{U}_{H}$ finishes a computation task ordered by $\mathcal{T}$, it performs blind decryption on the encrypted results and sends it back to $\mathcal{T}$ for local decryption.
Similar to many privacy-preserving computation schemes using HE[49, 6, 50], we employ the semi-honest threat model for the proposed protocol, wherein the server $\mathcal{U}_{H}$ is assumed to be a passive adversary that adheres to the protocol but still maintains curiosity about $\mathcal{T}$’s private data.
Note that verification of the outsourced decryption is required if we are to consider scenarios with higher security demands. However, we will not delve into this in the current paper, as verifying the correctness of the homomorphic computation itself is a non-trivial task and remains a significant research challenge today[51, 52, 53, 54]. Nevertheless, we will discuss feasible approaches to verify the correctness of the proposed protocol at the end of the paper for future research reference.
For our outsourcing decryption protocol to be applicable to the aforementioned scenario, it must adhere to the following properties[30]:
@@
Correctness For all arithmetic functions $f:\mathcal{M}^{I}\rightarrow\mathcal{M}$ defined in the underlying HE scheme, there exists an instance of the outsourcing decryption protocol such that $\Pi_{\text{LocDec}}(\Pi_{\text{BldDec}}(\text{HE.Eval}(f,pk,ct_{1},\dots,ct_{I}),\tilde{sk}),sk_{ob})=f(m_{1},\dots,m_{I})$ holds with overwhelming probability.
@@ IV-A Secret Key Blinding from Sparse Polynomial Inversion
From the description above, one can observe that routines $\Pi_{\text{SkBd}}$ and $\Pi_{\text{BldDec}}$ are just normal polynomial multiplications (see Algorithms 2 and 3). Note that the former one is invoked only once in the protocol, whose computational complexity can be amortized across multiple subsequent ciphertext decryption operations, making it negligible in the overall computational load. The latter is typically performed by $\mathcal{U}_{H}$ for each ciphertext to be decrypted.
The remaining challenge is to generate blinding factors that ensure $\Pi_{\text{LocDec}}$ operates with maximal efficiency. Note that the sparse-to-dense polynomial multiplication exhibits lower complexity compared to NTT-based dense multiplication when the sparsity and the circuit design are appropriately handled. Therefore, we choose sparse polynomials as the unblinding factors in $\Pi_{\text{SkBdKeyGen}}$. The sampling procedure is outlined in Algorithm 1, where we use hamming weight (denoted as $h$) to control the sparsity of a polynomial, and $S_{h}$ the set of all sparse polynomials over $R_{q}$ with hamming weight $h$. The unblinding factor is sampled from $S_{h}$ uniformly and randomly, with its inverse being the blinding factor. Note that an invertibility test should be done for the candidate unblinding factor since $R_{q}$ is merely a ring rather than a field.
@@
Elements individually satisfying the invertibility condition from different $R_{q_{i}}$ can be combined into an invertible element in $R_{q}$ using the CRT map (see Theorem IV.1). It can also be proven that elements sampled this way are uniformly distributed across $S_{h}$ (see Theorem IV.2). Using CRT can significantly reduce the complexity of the invertibility check since $q$ may be very large as we mentioned previously. Note that in homomorphic encryption schemes, the structure $q=\prod_{i=0}^{L}q_{i}$ is typically used to construct a large $q$, where $q_{i}$ are primes with similar length and satisfy the condition $q_{i}\equiv 1\mod{2d}$ to enable NTT-based polynomial multiplication.
Algorithm 1 Blinding key generation ($\Pi_{\text{SkBdKeyGen}}$)
0:  Ring $R_{q}=\mathbb{Z}_{q}[X]/X^{d}+1$; Hamming weight $h$
0:  Unblinding key $t$ with hamming weight $h$; Blinding key $t^{-1}\in R_{q}$
@@ IV-B Local Decryption
Algorithm 2 Secret Key Blinding ($\Pi_{\text{SkBd}}$)
0:  Blinding key $s\in R_{q}$; Secret key $s\in R_{q}$
0:  Blinded secret key $\tilde{s}\leftarrow ts\in R_{q}$
Algorithm 3 Blind Decryption ($\Pi_{\text{BldDec}}$)
0:  Blinded secret key $\tilde{s}$; Ciphertext $c=(u,v)\in R_{q}^{2}$
0:  Blinded decryption result $\tilde{c}\leftarrow(u\tilde{s},v)\in R_{q}^{2}$
Algorithm 4 Local Decryption ($\Pi_{\text{LocDec}}$)
0:  Unblinding key $t\in R_{q}$ with hamming weight $h$; Blinded decryption result $\tilde{c}=(\tilde{u},v)\in R_{q}$;
0:  Unblinded decryption result $m\in R_{q}$;
@@ IV-C Complexity Analysis
$\mathcal{T}$ is responsible for routine $\Pi_{\text{SkBdKeyGen}},\Pi_{\text{SkBd}}$ and $\Pi_{\text{LocDec}}$ in the protocol, where the former two are only performed once following the generation of the secret key, and the later is done for each ciphertext blindly decrypted by $\Pi_{\text{BldDec}}$. Therefore, the complexity of $\mathcal{T}$ is dominated by the sparse-to-dense polynomial multiplication with roughly $hd$ (modular) scalar multiplications.
On the other hand, the original decryption computed by $\mathcal{T}$ has a typical complexity of $ad\log{d}+d$ scalar multiplications corresponding to one Inversed NTT operation and coordinate-wise multiplication, where $a$ is a constant varies from concrete implementations of NTT. Then the efficiency of our proposed protocol is computed as $\frac{h}{a\log{d}+1}$.